OpenStudy (anonymous):

X^4+14x^2-32=0 Find all the imaginary solutions by factoring

OpenStudy (solomonzelman):

\[x^4+14x^2-32=0\]let x^2=a

OpenStudy (solomonzelman):

\[a^2+14a-32=0\]\[(a + 14)(a - 2 )=0\]\[a=-14~~~~~~and~~~~~~a=2\]

OpenStudy (solomonzelman):

\[x^2=a~~~~~~~~so\]\[x^2=-14~~~~~~~x=±i \sqrt{14}\]\[x^2=2~~~~~~~x=± \sqrt{2}\]

OpenStudy (anonymous):

Why would you let x^2=a

OpenStudy (anonymous):

\[x ^{4}+14x^2-32=0\]\[(x^2+16)(x^2-2)\] \[x^2=-16\]\[\sqrt{x}=\sqrt{-16}\]\[x=\pm4i\] \[x^2=2\]\[\sqrt{x}=\sqrt{2}\]\[x=\pm \sqrt{2}\]

OpenStudy (anonymous):

Oh thank you That makes more sense

OpenStudy (solomonzelman):

yes true it does, it become alg 1 from college alg. by simply saying let x^2=a

OpenStudy (anonymous):

no problem